unseen:

Solve for y. y - 12 = -10 y = 22 y = -22 y = -2 y = 2

unseen:

heeelp

KylieAurealis:

Add -12 to -10

unseen:

-22

unseen:

right?

KylieAurealis:

Itd be 2. If you were to do -10 - 12 then it would be -22. But you add -10 + 12

unseen:

no that y-12 is a -12 so you would add -12 +-10 so it would be -22

unseen:

my math teacher said that

AngeI:

y-12=-10 add 12 to both sides y-12+12=-10+12 getting y=2

AngeI:

No offense, your math teacher is wrong

KylieAurealis:



NOpanda14:

I agree with her

unseen:

ok i got it